OpenStudy (anonymous):

can you plz help as much as you can 1)Select the choice that translates the following verbal phrase correctly to an algebraic expression: the difference of c and 10 c + 10 10c 10 ÷ c c − 10 2)The official weight, in pounds, of four high school wrestlers is as follows: Wrestler A Wrestler B Wrestler C Wrestler D Weight 124.5 125.3 124.9 124.2 If the weight data are accurate to the nearest ones place only, which wrestler has a different weight? Wrestler A Wrestler B Wrestler C Wrestler D

OpenStudy (anonymous):

3)Evaluate the following expression using the values given: Find 2x − 3y − z if x = −2, y = 3, and z = −2 −11 11 −15 15 4)Martin wants to plot a graph to show the prices of different numbers of helmets. He has to plot the following points: (1, 24.97), (2, 49.94), (3, 74.91), (4, 99.88) Which of the following best describes how the y-axis should be labeled? From 0 to 300 at intervals of 50 From 0 to 100 at intervals of 25 From 0 to 500 at intervals of 100 From 0 to 80 at intervals of 10 5)It took Everly 55 minutes to run a 10-kilometer race last weekend. If you know that 1 kilometer equals 0.621 mile, how many minutes did it take Everly to run 1 mile during the race? Round your answer to the nearest hundredth. 3.41 minutes 5.50 minutes 6.21 minutes 8.86 minutes

OpenStudy (anonymous):

@Koikkara

OpenStudy (anonymous):

@Study_man can you help me

OpenStudy (koikkara):

@itsnotjesse the difference of c and 10 .... just the waqy you subtract \(2−3=-1\)

OpenStudy (anonymous):

so it would be d right because you subtract @Koikkara

OpenStudy (anonymous):

@Koikkara right?

OpenStudy (anonymous):

@Koikkara ?

OpenStudy (anonymous):

@Study_man can you help me plz

OpenStudy (koikkara):

@itsnotjesse yes \(c-10\)

OpenStudy (anonymous):

@Koikkara ok thank you and can you help me with any more ?

OpenStudy (koikkara):

these are the four different weights: Weight 124.5 125.3 124.9 124.2 @itsnotjesse check which one is not \(below~.5\)

OpenStudy (anonymous):

124.9 and 124.5 but i think its 124.9

OpenStudy (koikkara):

Here, if the decimal value is \(= or < than~ 5, ~the ~decimal~value~changes~i.e,, \) 124.9=125. \(you~are~right~!!\) @itsnotjesse

OpenStudy (anonymous):

@Koikkara ok thank you is that all or can you help with more

OpenStudy (koikkara):

next one, \(2x − 3y − z ,~x = −2, y = 3, ~z = −2\) substitute the values..

OpenStudy (anonymous):

im sorry can you help me a bit more to jog my memory i dint remember how to do this

OpenStudy (koikkara):

in place of x,y,z put the values, you are done

OpenStudy (koikkara):

\(2x-3y-z=0\) \(2*(-2)-3*(3)-(-2)=0\)

OpenStudy (anonymous):

so x,y,z i would replace them with -11 and see if it works and also with the others

OpenStudy (koikkara):

\(-4-9+2\)= \(-11\) @itsnotjesse

OpenStudy (anonymous):

so i was right the answer was -11

OpenStudy (anonymous):

?

OpenStudy (koikkara):

Yes

OpenStudy (anonymous):

ok thank you can you help me with the last two?

OpenStudy (koikkara):

@itsnotjesse next one... the point has the form of \((x,y)\) so there are 4 points given, can you tell me what is the maximum value of the y given ? ?

OpenStudy (anonymous):

99.88

OpenStudy (koikkara):

Correct, so we just need the y axis to be that long. so what interval of y you think we can take ?

OpenStudy (anonymous):

25?

OpenStudy (koikkara):

@itsnotjesse c'mon... you did everything by yourself yes..you are right

OpenStudy (anonymous):

ok so it's B

OpenStudy (koikkara):

Yes, and the last one, is here... http://openstudy.com/study#/updates/557f391ee4b07028ea60595b @itsnotjesse \(\color{blue}{\huge\tt{Nice~}}\) \(\huge\mathcal{to~meet~}\) \(\color{red}{\huge\tt{You~!}}\)

OpenStudy (anonymous):

@Koikkara ok thanks for all your help